# Aligning Speech to Languages to Enhance Code-switching Speech Recognition

### v2

@@ Abstract
Code-switching (CS) refers to the switching of languages within a speech signal and results in language confusion for automatic speech recognition (ASR). To address language confusion, we introduce a novel language alignment loss into ASR training to align acoustic features to pseudo-language labels learned from the ASR decoder. This approach enables frame-level language identification without the need for frame-level language annotations. To further tackle the complex token alternatives for language modeling in bilingual scenarios, we propose to employ large language models via a generative error correction method. A linguistic hint, which is derived from LAL outputs and decoded hypotheses, is introduced to guide the prompting and enhance the LLM-based generative error correction for CS-ASR. The proposed methods are evaluated on the SEAME dataset and data from the ASRU 2019 Mandarin-English code-switching speech recognition challenge. The incorporation of the proposed language alignment loss improves the CS-ASR performance for both hybrid CTC/attention and Whisper models on both datasets, with only a negligible increase in the number of parameters. This work also highlights the efficacy of language alignment loss in balancing primary-language-dominant bilingual data during training, with an 8.6% relative improvement on the ASRU dataset compared to the baseline model. Performance evaluation using large language models reveals the advantage of the linguistic hint by achieving 14.1% and 5.5% relative improvement on test sets of the ASRU and SEAME datasets, respectively.
@@ I Introduction
Although a CS-ASR system can function like monolingual ASR by combining language-specific vocabularies [1, 2], recent works address language confusion by incorporating language information. One direct approach is to optimize the ASR and language identification (LID) or diarization (LD) tasks jointly [3, 4, 5, 6, 7]. Here, models learn language information through backpropagation from the LID or LD branch during training [6, 8], while only the ASR output is computed during inference. Beyond joint optimization, approaches based on the bi-encoder and the mixture-of-experts method building upon the Transformer architecture [9, 10, 11, 12], where the models incorporate two encoders pre-trained on monolingual data independently to capture language-specific information. Language-specific modules have also been adopted in other architectures due to their effectiveness in distinguishing languages [13, 14]. In contrast to language-specific encoder modules, the language-aware decoder module has been explored to reduce multilingual contextual information via a language-specific self-attention mechanism within the Transformer decoders [15]. In addition, a conditional factorization method factorizes CS-ASR into two monolingual recognition processes before integrating multiple recognized monolingual segments into a single bilingual sequence [16]. As an extension, a conditionally factorized connectionist temporal classification (CTC) module that allows for a zero-shot setting has been proposed [17, 18]. A factored language model integrating syntactic and semantic features in a code-switched language model has also been explored to enhance CS-ASR [19].
@@
Although existing works achieved reasonable CS-ASR performance, constraints that impede model performance persist. Multilingual ASR systems often benefit from utterance-level one-hot language vectors, as each utterance typically contains only one language [20, 21]. In contrast, code-switching speech signals frequently contain two languages. Therefore, detecting language change points at high granularity (in particular at frame level) is preferable to utterance-level vectors [7]. Despite this need, ground-truth code-switching timestamps are often excluded in code-switching corpora since the human annotation process is resource-intensive and requires expertise in bilingualism. While force alignment offers a potential solution to generate high-granularity timestamps, code-switching speech poses greater challenges for this approach compared to monolingual speech due to increased language confusion [22, 23, 24]. Consequently, language labels generated through force alignment might not be desirable in terms of accuracy and computational cost. These limitations render the application of supervised language identification for joint optimization impractical [5]. Additionally, the accent of the primary language may bias the secondary language in bilingual code-switching speech, making the two languages auditorily similar [24, 25]. Therefore, achieving language identification in accented speech remains challenges [26].
@@
In this paper, we propose to leverage the speech-to-language alignment to improve CS-ASR performance. Central to our approach is introducing a language alignment loss (LAL), which enriches the CS-ASR model with language information without requiring additional annotations such as human labeling or force alignment. The LAL achieves this by utilizing low-granularity yet accurate token-level language information to explicitly guide high-granularity frame-level acoustic features. Unlike auxiliary frame-level LID methods, the proposed approach accounts for the proportion of tokens from each language and their auditory similarity, ensuring a more nuanced alignment between language and acoustic representations. Additionally, as the by-product of LAL, frame-level language predictions can be summarized as an utterance-level hint. This linguistic hint further utilizes the language information derived from speech to facilitate the incorporation of an external large-scale language model (LLM) in ASR via a generative error correction method [27, 28, 29].
@@
The remainder of this paper is organized as follows: Section II introduces the hybrid CTC/attention model, which serves as a baseline model on which we apply the proposed method. The proposed LAL and the incorporation of the code-switching hint in external language modeling are presented in Sections III and III-D, respectively. Information pertaining to the datasets, model configurations, and experiment setup are provided in Section IV. We present results and analysis in Section V before highlighting the advantages and limitations of our proposed methods in Section VI. Finally, we conclude our work in Section VII.
@@ II-A Conformer-based hybrid CTC/attention ASR model
We employ a Conformer-based hybrid CTC/attention ASR model as the baseline in this paper. The hybrid CTC/attention model comprises an encoder, a decoder, and a CTC module, where the decoder and CTC modules share the encoder outputs [30]. The encoder and decoder modules comprise Conformer encoder and Transformer decoder layers [31, 9], respectively [2].
Given a speech signal, we define its acoustic features as $\mathbf{X}=(\mathbf{x}_{t}\in\mathbb{R}^{F}|t=1,\ldots,T)$ and paired token sequence as $W=(w_{n}\in\mathcal{V}|n=1,\ldots,N)$, where $\mathcal{V}$ is the vocabulary, $T$ and $N$ are the lengths of feature and token sequences, respectively, and $F$ is the dimension of acoustic feature. The encoder generates hidden outputs $\mathbf{H}=(\mathbf{h}_{t}\in\mathbb{R}^{D}|t=1,\ldots,T^{\prime})$ from $\mathbf{X}$, which are subsequently used as inputs into the decoder and CTC modules. Here, $T^{\prime}$ is the length of the hidden output sequence, where $T^{\prime}<T$ due to the subsampling, and $D$ is the dimension of the hidden output. With $\mathbf{H}$, the CTC module computes the token sequences in accordance with the Bayesian decision theory by factorizing $p_{\mathrm{ctc}}\left(W|\mathbf{X}\right)$ as [18]
@@
is the acoustic model of CTC, where the probabilistic chain rule and the conditional independence assumption have been invoked. Exploiting the Bayes’ rule, the probabilistic chain rule, and the conditional independence assumption, the CTC token model is given by
@@
Consequently, the term $p\left(W\right)/p\left(Z\right)$ is often excluded and hence (1) can be rewritten as
@@
We note that tokens $W$ are embedded into $\mathbf{W}=(\mathbf{w}_{n}\in\mathbb{R}^{D}|n=1,\ldots,N)$ before being fed into the decoder module along with $\mathbf{H}$. The decoder then predicts the next token $w_{n}$ based on historical tokens $w_{1:n-1}$ and $\mathbf{H}$ via
@@
Here, $\widehat{W}$ is also referred to as a single hypothesis, and the final hypothesis of the given speech signal is chosen as the one with the highest likelihood among multiple hypotheses generated during the beam search. While the hybrid CTC/attention model has shown to be effective for the CS-ASR task, it performs CS-ASR similarly to monolingual ASR without exploiting any code-switching information, which, as a consequence, limits the CS-ASR performance.
Figure 1: Generative error correction method for improving ASR with an external LLM proposed in [29]. Variable $r$ is the low intrinsic rank, and matrices $\mathbf{A}$ and $\mathbf{B}$ are initialized by a random Gaussian distribution and zero, respectively. The red and black tokens within the N-best hypotheses denote the wrongly and correctly predicted tokens, respectively.
@@ II-B Improving ASR via LLM and efficient fine-tuning
Due to the scarcity of code-switching data, general code-switching language models moderately improve the performance of a CS-ASR system when being incorporated through shallow fusion [4, 7]. Although developing an external language model on monolingual data or synthesized code-switching data has shown to be effective [33, 34, 35], the performance is limited by the domain mismatch. Since large language models have achieved success in natural language processing and have been extended for computer vision and speech signal processing applications [36, 27, 37], we propose to adopt open-source LLMs, which is robust against diverse domains due to the large-scale training data, to improve CS-ASR by addressing the complex token alternatives in bilingual scenarios.
Recent works have also attempted to improve speech recognition through the use of LLMs. A direct method involves prompting an LLM using paired discrete speech and text embeddings [38]. A generative error correction method has also been applied to LLMs [29] as presented in Fig. 1, where the final prediction is generated by summarizing and correcting the N-best ASR hypotheses [39]. This approach has shown to be effective in monolingual ASR. However, code-switching leads to more token alternatives that are of similar auditory or syntactic characteristics compared to a monolingual application—direct transference of the generative error correction to CS-ASR may not be desirable.
@@
where $\mathbf{\Delta M}$ is the model parameters of the model update. Matrices $\mathbf{A}$ and $\mathbf{B}$ are initialized as by random Gaussian and zero values, respectively, so that $\mathbf{\Delta M}=\mathbf{AB}=0$ at the beginning of training.
@@ III Language Alignment Loss
Since multilingual ASR benefits from the supplementary language information offered by utterance-level one-hot language vectors, we propose to enhance CS-ASR performance by incorporating frame-level language information, enabling the detection of code-switching at high granularity. In this work, we introduce the LAL that is incorporated to the encoder-decoder framework to capture language information as shown in Fig. 2.
@@ III-A Frame-level language identification
To capture frame-level language information, we employ a linear layer as a built-in language classifier. This layer takes the hidden output unit of the encoder module as its input and aims to generate a language decision for each hidden output unit.
@@
Although frame-level ground-truth language timestamps are unavailable, token-level language information can readily be inferred from text, particularly in cases where there is a notable contrast in character structure or morphology between the two languages. Past research has investigated the conversion of byte-pair encoding (BPE) tokens into their respective language labels to facilitate language identification or diarization [41, 4, 42]. Here, we employ a similar conversion strategy, where sub-tokens like BPE tokens are first transformed into token-level labels corresponding to their respective languages. The pseudo-frame-level language annotations are subsequently extracted by aligning frames to these token-level language labels.
@@ III-B Aligning frames with token-level language labels
The encoder-decoder model achieves ASR by mapping speech features to tokens. The alignment between speech frames (i.e., encoder hidden outputs) and text (i.e., tokens) is inherently learned through the cross-attention process illustrated on the left of Fig. 3. Given multiple attention heads within the last Transformer decoder layers, their cross-attention matrices are averaged to generate the weight matrix $\mathbf{Atten}\in\mathbb{R}^{T^{\prime}\times N}$ of the speech-to-text alignment shown in the top matrix of Fig. 3. The alignment between hidden outputs and languages can then be derived from the frame-to-token weight matrix illustrated by the transition from the top to bottom matrices in Fig 3. Specifically, the averaged cross-attention weight matrix can be decomposed into vectors along $T^{\prime}$, being $\mathbf{Atten}=(\mathbf{atten}_{t}\in\mathbb{R}^{N}|t=1,\ldots,T^{\prime})$. Each element in $\mathbf{atten}_{t}$ denotes the attention weight corresponding to a BPE token within the input token sequence. Each frame can then be assigned a pseudo language label which is the language of the BPE token corresponding to the highest weight. This is achieved via
@@
[…]
The pseudo language labels are concatenated to form a sequence $\mathbf{Y}=(\mathbf{y}_{t}\in\mathbb{R}^{C}|t=1,\ldots,T^{\prime})$. Here, $T^{\prime}$ is the number of hidden output units and $C$ is the number of languages, where all special tokens are treated as a third language in addition to the two target languages. Each hidden output unit is subsequently projected to a language decision $\mathbf{y}_{t}$ by a language classifier comprising one linear layer
@@
where $\mathbf{\widehat{y}}_{t,i}$ denotes the value at the $i$th dimension of the language prediction vector for the $t$th hidden output vector. Similarly, $\mathbf{y}_{t,c}$ denotes the value at the $c$th dimension of the one-hot pseudo language label vector for the $t$th hidden output vector such that $\mathbf{y}_{t,c}=1$ for the target language with the remaining elements being zero.
The key distinction between the proposed LAL and general LID lies in the ability of the proposed LAL to iteratively refine pseudo-labels during training, enabling a mutual optimization process between ASR and LID tasks. The performance of the CS-ASR model improves during model training as the language information is incorporated, reaching its peak upon being optimized completely. The heightened ASR performance contributes to increasingly accurate language labels. This improvement, in turn, leads to higher language identification performance. Therefore, extracting pseudo language labels and language identification has become integral to the iterative optimization process during training.
Figure 3: The pseudo frame-to-language label conversion process. On the right side, the upper matrix represents the cross-attention weight matrix averaged over all attention heads within the last decoder layer. The middle matrix represents the converted frame-to-language weight matrix and the pseudo language labels are illustrated at the bottom. In this illustrative example, red and green cells denote the English and Mandarin tokens, respectively.
III-C Balancing training via toke-level language weights

### v3

@@ Abstract
Code-switching (CS) refers to the switching of languages within a speech signal and results in language confusion for automatic speech recognition (ASR). To address language confusion, we propose a language alignment loss (LAL) that aligns acoustic features to pseudo-language labels learned from the ASR decoder during ASR training. This approach enables frame-level language identification without the need for frame-level language annotations. To further tackle the complex token alternatives for language modeling in bilingual scenarios, we propose to employ large language models via a generative error correction method. A linguistic hint, derived from LAL outputs and decoded hypotheses, is introduced to guide the prompting and enhance the LLM-based generative error correction for CS-ASR. The proposed methods are evaluated on the SEAME dataset and data from the ASRU 2019 Mandarin-English code-switching speech recognition challenge. The incorporation of the proposed language alignment loss improves CS-ASR performance for both hybrid CTC/attention and Whisper models on both datasets, with only a negligible increase in the number of parameters. This work also highlights the efficacy of language alignment loss in balancing primary-language-dominant bilingual data during training, with an 8.6% relative improvement on the ASRU dataset compared to the baseline model. Performance evaluation using large language models reveals the advantage of the linguistic hint by achieving 14.1% and 5.5% relative improvement on test sets of the ASRU and SEAME datasets, respectively.
@@ I Introduction
Although a CS-ASR system can function like monolingual ASR by combining language-specific vocabularies [1, 2, 3], recent works address language confusion by incorporating language information. One direct approach is to optimize the ASR and language identification (LID) or diarization (LD) tasks jointly [4, 5, 6, 7, 8, 9]. Here, models learn language information through backpropagation from the LID or LD branch during training [7, 10], while only the ASR output is computed during inference. Beyond joint optimization, approaches based on the bi-encoder and the mixture-of-experts method build upon the Transformer architecture [11, 12, 13, 14], where the models incorporate two encoders pre-trained on monolingual data independently to capture language-specific information. Language-specific modules have also been adopted in other architectures due to their effectiveness in distinguishing languages [15, 16]. In contrast to language-specific encoder modules, the language-aware decoder module has been explored to reduce multilingual contextual information via a language-specific self-attention mechanism within Transformer decoders [17]. In addition, a conditional factorization method factorizes CS-ASR into two monolingual recognition processes before integrating multiple recognized monolingual segments into a single bilingual sequence [18]. As an extension, a conditionally factorized connectionist temporal classification (CTC) module that allows for a zero-shot setting has been proposed [19, 20]. A factored language model integrating syntactic and semantic features in a code-switched language model has also been explored to enhance CS-ASR [21].
@@
Although existing works achieve reasonable CS-ASR performance, limitations in code-switching annotations and data characteristics continue to hinder further improvements. Multilingual ASR systems often benefit from utterance-level one-hot language vectors since each utterance typically contains only one language [22, 23]. In contrast, code-switching speech signals often contain two languages, making utterance-level language labels insufficient. To address this, detecting languages at finer granularity, particularly at the frame or token level, has been suggested as a more suitable approach to CS-ASR [4, 5, 6, 8, 9, 13, 16]. Despite this, most code-switching corpora do not include ground-truth frame-level language timestamps, since the human annotation process is resource-intensive and requires expertise in bilingualism. While forced alignment offers a potential solution to generate high-granularity timestamps, code-switching speech poses greater challenges for this approach compared to monolingual speech due to increased language confusion [24, 25, 26]. Consequently, language labels generated through forced alignment might not be desirable in terms of accuracy and computational cost. These limitations render the application of supervised language identification for CS-ASR impractical [6]. Additionally, the accent of the primary language may bias the secondary language in bilingual code-switching speech, resulting in the two languages being auditorily similar [26, 27]. Therefore, achieving language identification in accented speech remains challenging [28].
@@
In this paper, we propose leveraging the speech-to-language alignment to improve CS-ASR performance. Central to our approach is introducing a language alignment loss (LAL), which enriches the CS-ASR model with language information without the need for additional annotations such as human labeling or forced alignment. The LAL achieves this by utilizing low-granularity yet accurate token-level language information to explicitly guide high-granularity frame-level acoustic features. The proposed LAL relies on frame-level language information and is therefore categorized as a joint optimization approach. In addition to not requiring frame-level language annotations, the proposed method also differs from auxiliary LID approaches by incorporating token-level language proportions and auditory similarity, leading to a more nuanced alignment between language and acoustic representations. Additionally, as a by-product of LAL, frame-level language predictions can be summarized as an utterance-level hint. This linguistic hint further utilizes the language information derived from speech by facilitating the incorporation of an external large-scale language model (LLM) in ASR through a generative error correction method [29, 30, 31].
@@
The remainder of this paper is organized as follows: Section II introduces the hybrid CTC/attention model, which serves as a baseline model for the proposed method. The proposed LAL and the incorporation of the code-switching hint in external language modeling are presented in Sections III and III-D, respectively. Datasets, model configurations, and experimental setup are described in Section IV. We present results and analysis in Section V before highlighting the advantages and limitations of our proposed methods in Section VI. Finally, we conclude our work in Section VII.
@@ II-A Conformer-based hybrid CTC/attention ASR model
We employ a Conformer-based hybrid CTC/attention ASR model as the baseline. The hybrid CTC/attention model comprises an encoder, a decoder, and a CTC module, where the decoder and CTC modules share the encoder outputs [32]. The encoder and decoder modules comprise the Conformer encoder and Transformer decoder layers [33, 11, 2], respectively.
Given a speech signal, we define its acoustic features as $\mathbf{X}=(\mathbf{x}_{t}\in\mathbb{R}^{F}|t=1,\ldots,T)$ and the paired token sequence as $W=(w_{n}\in\mathcal{V}|n=1,\ldots,N)$, where $\mathcal{V}$ is the vocabulary, $T$ and $N$ are the lengths of the feature and token sequences, respectively, and $F$ is the dimension of the acoustic feature. The encoder generates hidden outputs $\mathbf{H}=(\mathbf{h}_{t}\in\mathbb{R}^{D}|t=1,\ldots,T^{\prime})$ from $\mathbf{X}$, which are subsequently used as inputs for the decoder and CTC modules. Here, $T^{\prime}$ is the length of the hidden output sequence, where $T^{\prime}<T$ due to subsampling, and $D$ is the dimension of the hidden output. With $\mathbf{H}$, the CTC module computes the token sequences according to the Bayesian decision theory by factorizing $p_{\mathrm{ctc}}\left(W|\mathbf{X}\right)$ as [20]
@@
is the acoustic model of the CTC, where the probabilistic chain rule and the conditional independence assumption have been invoked. Exploiting the Bayes’ rule, the probabilistic chain rule, and the conditional independence assumption, the CTC token model is given by
@@
The scaling term $p\left(W\right)/p\left(Z\right)$ is often excluded, and hence (1) can be rewritten as
@@
We note that the tokens $W$ are embedded into $\mathbf{W}=(\mathbf{w}_{n}\in\mathbb{R}^{D}|n=1,\ldots,N)$ before being fed into the decoder module along with $\mathbf{H}$. The decoder then predicts the next token $w_{n}$ based on historical tokens $w_{1:n-1}$ and $\mathbf{H}$ via
@@
Here, $\widehat{W}$ is also referred to as a single hypothesis, and the final hypothesis of the given speech signal is chosen as the one with the highest likelihood among multiple hypotheses generated during the beam search. While the hybrid CTC/attention model has proven to be effective for the CS-ASR task, it performs CS-ASR similarly to monolingual ASR without exploiting any code-switching information, which consequently limits the CS-ASR performance.
Figure 1: Generative error correction method for improving ASR with an external LLM proposed in [31]. Variable $r$ is the low intrinsic rank, and matrices $\mathbf{A}$ and $\mathbf{B}$ are initialized from a zero-mean Gaussian distribution and as zeros, respectively. The red and black tokens within the N-best hypotheses denote the wrongly and correctly predicted tokens, respectively.
@@ II-B Improving ASR via LLM and efficient fine-tuning
Due to the scarcity of code-switching data, general code-switching language models moderately improve the performance of a CS-ASR system when incorporated through shallow fusion [5, 8]. Although developing an external language model on monolingual data or synthesized code-switching data is effective [35, 36, 37], the performance is limited by domain mismatch. Since large language models have achieved success in natural language processing and have been extended to computer vision and speech signal processing applications [38, 29, 39], we propose to adopt open-source LLMs, which are robust against diverse domains due to the large-scale training data, to improve CS-ASR by addressing complex token alternatives in bilingual scenarios.
Recent works have also attempted to improve speech recognition through the use of LLMs. A direct method involves prompting an LLM using paired discrete speech and text embeddings [40]. A generative error correction method has also been applied to LLMs [31] as illustrated in Fig. 1, where the final prediction is generated by summarizing and correcting the N-best ASR hypotheses [41]. This approach has shown effectiveness in monolingual ASR. However, code-switching leads to more token alternatives that have similar auditory or syntactic characteristics compared to a monolingual application—direct transference of the generative error correction to CS-ASR may not be desirable.
@@
where $\mathbf{\Delta M}$ is the model parameters of the model update. Matrices $\mathbf{A}$ and $\mathbf{B}$ are initialized from a zero-mean Gaussian distribution and as zeros, respectively, such that $\mathbf{\Delta M}=\mathbf{AB}=0$ before training.
@@ III Language Alignment Loss
Since multilingual ASR benefits from the supplementary language information offered by utterance-level one-hot language vectors, we propose to enhance CS-ASR performance by incorporating frame-level language information, enabling the detection of code-switching at high granularity. To this end, and as shown in Fig. 2, we introduce the LAL that is incorporated into the encoder-decoder framework for capturing language information.
@@ III-A Frame-level language identification
To capture frame-level language information, we employ a linear layer as a built-in language classifier. This layer takes the hidden output units of the encoder module as its input and aims to generate a language decision for each hidden output unit.
@@
Although frame-level ground-truth language timestamps are unavailable, token-level language information can readily be inferred from text, particularly in cases where there is a notable contrast in character structure or morphology between the two languages. Past research has investigated the conversion of byte-pair encoding (BPE) tokens into their respective language labels to facilitate language identification or diarization [43, 5, 44]. Here, we employ a similar conversion strategy, where sub-tokens such as the BPE tokens are first transformed into token-level labels corresponding to their respective languages. The pseudo-frame-level language annotations are subsequently extracted by aligning frames to these token-level language labels.
@@ III-B Aligning frames with token-level language labels
The encoder-decoder model achieves ASR by mapping speech features to tokens. The alignment between speech frames (i.e., encoder hidden outputs) and text (i.e., tokens) is inherently learned through the cross-attention process illustrated on the left of Fig. 3. Given multiple attention heads within the last Transformer decoder layers, their cross-attention matrices are averaged to generate the weight matrix $\mathbf{Atten}\in\mathbb{R}^{T^{\prime}\times N}$ of the speech-to-text alignment shown in the top matrix of Fig. 3. The alignment between hidden outputs and languages can then be derived from the frame-to-token weight matrix as illustrated from the top to bottom matrices in Fig 3. Specifically, the averaged cross-attention weight matrix can be decomposed into vectors along $T^{\prime}$, being $\mathbf{Atten}=(\mathbf{atten}_{t}\in\mathbb{R}^{N}|t=1,\ldots,T^{\prime})$. Each element in $\mathbf{atten}_{t}$ denotes the attention weight corresponding to a BPE token within the input token sequence. Each frame can then be assigned a pseudo-language label, which is the language of the BPE token corresponding to the highest weight. This is achieved via
@@
where $\mathrm{T2L}(\cdot)$ represents the conversion from the $n$-th BPE token with the highest attention weight in $\mathbf{atten}_{t}$ to its language label.
The pseudo-language labels are concatenated to form a sequence $\mathbf{Y}=(\mathbf{y}_{t}\in\mathbb{R}^{C}|t=1,\ldots,T^{\prime})$. Here, $T^{\prime}$ is the number of hidden output units and $C$ is the number of languages, where all special tokens are treated as a third language in addition to the two target languages. Each hidden output unit is subsequently projected to a language decision $\mathbf{y}_{t}$ by a language classifier comprising one linear layer
@@
where $\mathbf{\widehat{y}}_{t,i}$ denotes the value at the $i$th dimension of the language prediction vector for the $t$th hidden output vector. Similarly, $\mathbf{y}_{t,c}$ denotes the value at the $c$th dimension of the one-hot pseudo-language label vector for the $t$th hidden output vector such that $\mathbf{y}_{t,c}=1$ for the target language with the remaining elements being zero. The proposed LAL is employed on the encoder representations, enriching the ASR encoder (acoustic end) with frame-level language information. The ASR decoder is not explicitly supervised by this objective function.
The key distinction between the proposed LAL and general LID lies in the ability of the former to refine pseudo-labels iteratively during training, enabling a mutual optimization process between ASR and LID tasks. The performance of the CS-ASR model improves during model training as the language information is incorporated, reaching its peak performance upon being optimized completely. The improved ASR performance contributes to increasingly accurate language labels. This improvement, in turn, leads to higher language identification performance. Therefore, extracting pseudo-language labels and language identification has become integral to the iterative optimization process during training.
Figure 3: The pseudo frame-to-language label conversion process. On the right side, the upper matrix represents the cross-attention weight matrix averaged over all attention heads within the last decoder layer. The middle matrix represents the converted frame-to-language weight matrix and the pseudo-language labels are illustrated at the bottom. In this illustrative example, red and green cells denote the English and Mandarin tokens, respectively.
III-C Balancing training via token-level language weights